OpenStudy (anonymous):

suppose g(s)=(s^2)+3s+1. find a value c in the interval [0,5] such that g'(c) equals the average rate of change of g(s) on the interval [0,5].

OpenStudy (anonymous):

im getting -3/2 or -1.5

OpenStudy (jamesj):

What is the average rate of change of g on that interval?

OpenStudy (anonymous):

41/5

OpenStudy (jamesj):

No

OpenStudy (jamesj):

(g(5) - g(0))/(5-0) = ( 41 - 1 ) / 5 = 8

OpenStudy (anonymous):

its plus 1

OpenStudy (jamesj):

Yes, and g(0) = 1. Now, what's the derivative, g'(s)? And the solution of g'(s) = 8?

OpenStudy (anonymous):

o nvm

OpenStudy (anonymous):

2s+3

OpenStudy (anonymous):

@JamesJ

OpenStudy (jamesj):

Now you want a value of s such that g'(s) = 8 i.e., 2s + 3 = 8

OpenStudy (anonymous):

its 2.5. thanks.

OpenStudy (jamesj):

yes